Suppose the Japanese yen exchange rate is ¥70.47 = $1, and the British pound exchange rate is £1 = $1.50.

a.
What is the cross-rate in terms of yen per pound? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16. Do not include the yen sign (¥).)

Cross-rate            ¥

b. Suppose the cross-rate is ¥109 = £1. What is the arbitrage profit per dollar? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., 32.1616.)

Arbitrage profit per dollar            $
Answer #1

Cross Rate of yen per pound = (Yen/Dollar)*(Dollar/Pound)

= (70.47)*1.50

= 105.705 Yen/Pound

b.Actual Rate is Yen 109/Pound

Convert 1 Dollar into pound and get 1/1.5 = 0.66667

Convert Pound into Yen and Get 0.666667*109 = Yen 72.666667

Convert Yen into Dollar = 72.67/70.47 = $1.03117

Hence, arbitrage profit per dollar = 1.03117-1

= $0.03117

i.e. 0.0312